OpenStudy (anonymous):

A store has a total of 15 bicycles and tricycles. There are 37 wheels. How many tricycles does the store have?

OpenStudy (kropot72):

Let the number of bicycles be b and the number of tricycles be t. Then we can write the equation: b + t = 15 ................(1) We can also write another equation based on the number of wheels: 2b + 3t = 37 ............(2) Equations (1) and (2) are a pair of simultaneous equations that can be solved to find the value of t, which is the number of tricycles. Can you solve for t?

OpenStudy (kropot72):

Equation (1) can be rearranged to give b in terms of t as follows: b + t = 15 ................(1) Subtracting t from both sides of (1) gives: b = 15 - t .................(3) If we plug (15 - t) into equation (2) instead of b we get: 2(15 - t) + 3t = 37 which can be expanded to give: 30 -2t +3t = 37 ..........(4) Now you can solve equation (4) to find the value of t.